EvieSwan2405:

Math! Help! Quiz: Graphing Linear Equations Which linear equation is represented in the graph? y = 2x – 1 y = –2x – 1 y = –2x 1 y = 2x 1 ____________________________ Also 5 questions!

pokie:

is there a graph

EvieSwan2405:

1 attachment
EvieSwan2405:

@Ultrilliam

EvieSwan2405:

This is what I came to ask

Felicity:

I thinks its D Evie

Felicity:

@EvieSwan2405

vaporeon246:

No it's B, second choice

sillybilly123:



sillybilly123:

In slope/intercept form, Americans say: \(y = mx + b\) the intercept, ie \(y(x=0)\), is: \(b = 1\) the slope is: \(m = \dfrac{\text{change in y}}{\text{change in x}} = \dfrac{-3-1}{2-0} =\dfrac{-4}{2} = -2\) so you have: \(y = -2 x + 1\) then check that. \(y(2) = -2(2) + 1 = -3\) \(y(0) = -2(0) + 1 = 1\) looks OK